Session: Feb 2001 (session 1)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Mauro at Tue Feb 13 04:06:01 MET 2001) Success
Any scan is recorded. Sky is clear, temp. = 6 C , gps- formatter = -84,7 uSec.

Torun (K.Borkowski at Tue Feb 13 10:10:44 MET 2001) Minor failures
During the scan from 2:37 antenna was off source due to long pass in the azimuth. No other known problems.

Effelsberg (A. Kraus at Tue Feb 13 11:00:44 MET 2001) Minor failures
The scans between UT0124-0135 were lost: problems with the antenna control computer.

Jodrell_Bank (Alastair Gunn at Tue Feb 13 17:37:57 MET 2001) Success
No known problems.

Shanghai (Huang Xinyong at Wed Feb 14 00:29:25 MET 2001) Success
The observing is recorded with VLBA formater.

Urumqi (Liu Xiang at Wed Feb 14 11:08:15 MET 2001) Success

Noto (C.Trigilio at Thu Feb 15 18:45:03 MET 2001) Success
No problem reported.

Westerbork (Tony Foley at Fri Feb 16 16:28:25 MET 2001) Minor failures
RT0 not working. First and last BBCs used rest of the array (RT1-D) and 2nd BBC used west (RT1-6). 3rd used east (RT7-D).

Westerbork (Tony Foley at Fri Feb 16 16:30:58 MET 2001) Success
Sorry -wrong experiment feedback last time. No problems.